# supplement

### v1

@@ 4 Theoretical Properties
where $r\geq 1$ is a fixed constant, and the weighted $\ell_{1}$ norm $\|\cdot\|_{1,w}$ and weighted $\ell_{2}$ norm $\|\cdot\|_{2,w}$ for a vector $a\in\mathbb{R}^{M}$ are given by $\|a\|_{1,w}:=\frac{1}{N}\sum_{m=1}^{M}n_{m}|a_{m}|$ and $\|a\|_{2,w}:=\sqrt{\frac{1}{N}\sum_{m=1}^{M}n_{m}a_{m}^{2}}$. Therefore, (17)-(18) quantify the entry-wise integrative deviation between $A^{(\cdot)}$ and $B^{(\cdot)}$ under the $r$-th power of the weighted $\ell_{1}$ and $\ell_{2}$ norms, respectively. It is noteworthy that the sample size heterogeneity across the local sites is fully accommodated by the weights $\{n_{m}/N\}_{m\in[M]}$, and consequently there is no necessity to impose the commonly assumed balanced sample size condition to derive the convergence rates.
@@ 4.1 Rates of Convergence for HEAT
The thresholding function $T_{1,\lambda}:\mathbb{R}\rightarrow\mathbb{R}$ satisfies the properties (15), and the function $T_{2,\lambda}:\mathbb{R}^{M}\rightarrow\mathbb{R}^{M}$ satisfies the properties (16) with $\|\cdot\|=\|\cdot\|_{2,w}$ on $\mathbb{R}^{M}$.
@@
Assumptions 1-3 impose mild conditions on data distributions, and Assumption 4 summarizes the requirements for the thresholding functions. Specifically, Assumption 1 is a regularity condition to prevent singular covariances and avoid diverging variances. Assumption 2 excludes the heavy-tailed distributions, and Assumption 3 assumes non-degenerate variances for the influence functions $\phi_{i,j,k}^{(m)}$’s. Assumptions 2-3 can be implied by Assumption 1 if the distributions are Gaussian. The weighted norm $\|\cdot\|_{2,w}$ for $T_{2,\lambda}$ in Assumption 4 accommodates the sample size differences among the distributed sites. Section LABEL:sec:choice_threshold_fun of the supplement presents specific examples of $T_{1,\lambda}$ and $T_{2,\lambda}$ that satisfy Assumption 4, including soft, hard, SCAD and MCP thresholding functions.
@@
Under Assumptions 1-2, assume $s_{0}\log(p)=o(n)$, $n\geq Cs_{0}^{2}\log(p)$ for some constant $C>0$, and $\lambda_{j}^{(m)}\asymp\sqrt{\log(p)/n_{m}}$ uniformly for $j\in[p]$ and $m\in[M]$, and define
@@
and $\tau=\widebar{C}_{3}M^{5/2}\log(p\vee n_{\max})\log(p)/n^{1/2}$ for some constants $\widebar{C}_{2},\widebar{C}_{3}>0$, then for any constant $\widebar{\delta}>0$, the event $\bigcap_{j,k\in[p]}\left\{\|\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}-\Lambda_{j,k}^{(\cdot)}\|_{2,w}\leq\widebar{\lambda}_{2,j,k}\right\}$ holds with probability tending to 1.
Propositions 1-2 characterize the volatility of $\widebar{\Omega}_{j,k}$ and $\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}$. As a result, $|\widebar{\Omega}_{j,k}|\leq\widebar{\lambda}_{1,j,k}$ and $\|\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}\|_{2,w}\leq\widebar{\lambda}_{2,j,k}$ respectively indicate that signals $\Gamma_{j,k}$ and $\Lambda_{j,k}^{(\cdot)}$ are weak, allowing for thresholding $\widebar{\Omega}_{j,k}$ or $\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}$ to exactly zero. The variance terms $\|v_{j,k}^{(\cdot)}\|_{1,w}$, $\|v_{j,k}^{(\cdot)}\|_{1},\|v_{j,k}^{(\cdot)}\|_{2},\|v_{j,k}^{(\cdot)}\|_{\infty}$ in $\widebar{\lambda}_{1,j,k}$ and $\widebar{\lambda}_{2,j,k}$ reflect entry-dependent volatility. Although these terms can be replaced by a universal constant for universal shrinkage levels, the proposed thresholds adapt to heteroscedasticity and provide practical guidance on shrinkage level selection. The remaining terms, $s_{0}M\log(p)/N$ in Proposition 1 and $s_{0}\sqrt{M}\log(p)/\sqrt{nN}$ in Proposition 2, represent higher-order errors of the local estimators (6). The additional term $\tau/\sqrt{N}$ in Proposition 2 is technical and potentially subject to further theoretical refinement. It arises from the Chi-square approximation for the sum-of-square-type quantity $\|\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}-\Lambda_{j,k}^{(\cdot)}\|_{2,w}$ as presented in Lemma LABEL:lemma:ss_chis_bound of the supplement. It is negligible if $M^{4}\log^{4}(p\vee n_{\max})=o(n)$, a simplification assumed in subsequent analyses for clarity. It’s essential to highlight that compared to typical concentration results for quadratic forms or polynomials, the Chi-square approximation developed here offers tight bounds on both constant terms and convergence rates. Therefore, it facilitates both practical implementation and theoretical error analysis.
In practice, the variances are unknown and we can replace them by the empirical estimator $\widehat{v}_{j,k}^{(m)}$ proposed in (9) of Section 3.1. The following Proposition 3 establishes the uniform rate of convergence for the variance estimation.
@@
By Propositions 1-3, for a small constant $\delta>0$, the shrinkage levels $\lambda_{1,j,k}$ and $\lambda_{2,j,k}$ in (13)-(14) can be determined by
@@
$\lambda_{2,j,k}=(1+\delta)\sqrt{\frac{\|\widehat{v}_{j,k}^{(\cdot)}\|_{1}+2\sqrt{2}\|\widehat{v}_{j,k}^{(\cdot)}\|_{2}\sqrt{\log(p)}+4\|\widehat{v}_{j,k}^{(\cdot)}\|_{\infty}\log(p)}{N}}+\widebar{C}_{2}\frac{\sqrt{M}s_{0}\log(p)}{\sqrt{nN}}.$ (20)
Subsequently, the following theorems establish the rates of convergence for the HEAT estimator $\widetilde{\Omega}^{(\cdot)}$ under the loss matrices $L_{1,r}$ and $L_{2,r}$ in (17)-(18). Specifically, Theorem 1 evaluates the integrative estimation losses under the general matrix $l$-norm with $l\in[1,\infty]$, while Theorem 2 investigates the matrix Frobenius norm, which essentially vectorizes the loss matrices and measures the accuracy through vector $\ell_{2}$ norm.
@@
Under Assumptions 1-4, if $n\geq Cs_{0}^{2}\log(p)$ for some constant $C>0$, $\log^{2}(p\vee n_{\max})(s_{0}\vee\log(p))=o(n)$, $M^{4}\log^{4}(p\vee n_{\max})=o(n)$, and $\lambda_{j}^{(m)}\asymp\sqrt{\log(p)/n_{m}}$ uniformly for $j\in[p]$ and $m\in[M]$, then for any $r\geq 1$ and $l\in[1,\infty]$, we have
@@
The convergence rates in Theorem 1 consist of two components: $s_{1}(\varphi_{1})^{r}$ and $s_{2}(\varphi_{2})^{r}$, which respectively reflect the estimation errors under the homogeneity and the heterogeneity. The same decomposition and interpretation also apply to Theorem 2. Note that both $\varphi_{1}$ and $\varphi_{2}$ consist of two terms. The terms $\sqrt{\frac{\log(p)}{N}}$ and $\sqrt{\frac{M+\log(p)}{N}}$ are the leading statistical errors for integration, while $\frac{s_{0}M\log(p)}{N}$ and $\frac{\sqrt{M}s_{0}\log(p)}{\sqrt{nN}}$ are the higher-order errors induced by individual estimation, representing the additional costs of distributed data analysis. Consequently, the higher-order errors will be dominated by the leading error terms, i.e.,
$\varphi_{1}\asymp\psi_{1}:=\sqrt{\frac{\log(p)}{N}}\text{ and }\varphi_{2}\asymp\psi_{2}:=\sqrt{\frac{M+\log(p)}{N}},$
if $\frac{M^{2}}{N}\lesssim{(s_{0}^{2}\log(p))}^{-1}$ and $s_{0}^{2}\log^{2}(p)\lesssim(1\vee\frac{\log(p)}{M})n$ hold. These restrictions imply that the number of distributed sites can not be too large, and similar requirements also appear in the distributed algorithms under homogeneity (e.g., Zhang et al.,, 2015; Lee et al.,, 2017; Battey et al.,, 2018; Jordan et al.,, 2019) and heterogeneity (e.g., Maity et al.,, 2022; Cai et al.,, 2022). In the following section, it will be shown that $\psi_{1}$ and $\psi_{2}$ achieve the minimax optimal rates.
@@ 4.2 Minimax Lower Bounds for Integrative Estimation
Suppose that $X_{i}^{(m)}\sim N(\mu^{(m)},\Sigma^{(m)})$ for $i\in[n_{m}],m\in[M]$ and Assumption 1 holds. If $23\leq s_{1}<p/$ 2, $22\leq s_{2}<p/$ 2, $s_{0}^{2}\log^{3}(p)=o(N)$, $s_{2}^{2}M=o(N)$, $N\lesssim nM$ and $p\gtrsim N^{\beta}$ for some $\beta>1$, then for any $r\geq 1$ and $l\in[1,\infty]$,
$\inf_{\widetilde{\Omega}^{(\cdot)}}\sup_{\Omega^{(\cdot)}\in\Theta}\mathbb{E}\|L_{1,r}(\widetilde{\Omega}^{(\cdot)},\Omega^{(\cdot)})\|_{l}\gtrsim s_{1}\psi_{1}^{r}+s_{2}\psi_{2}^{r},$
$\inf_{\widetilde{\Omega}^{(\cdot)}}\sup_{\Omega^{(\cdot)}\in\Theta}\mathbb{E}\|L_{2,r}(\widetilde{\Omega}^{(\cdot)},\Omega^{(\cdot)})\|_{l}\gtrsim s_{1}\psi_{1}^{r}+s_{2}\psi_{2}^{r}.$
@@
$\inf_{\widetilde{\Omega}^{(\cdot)}}\sup_{\Omega^{(\cdot)}\in\Theta}\frac{1}{p}\mathbb{E}\|L_{1,r}(\widetilde{\Omega}^{(\cdot)},\Omega^{(\cdot)})\|_{\text{F}}^{2}\gtrsim s_{1}\psi_{1}^{2r}+s_{2}\psi_{2}^{2r},$
$\inf_{\widetilde{\Omega}^{(\cdot)}}\sup_{\Omega^{(\cdot)}\in\Theta}\frac{1}{p}\mathbb{E}\|L_{2,r}(\widetilde{\Omega}^{(\cdot)},\Omega^{(\cdot)})\|_{\text{F}}^{2}\gtrsim s_{1}\psi_{1}^{2r}+s_{2}\psi_{2}^{2r}.$
Theorems 3-4 establish the minimax lower bounds under the matrix $l$-norm and Frobenius norm, respectively. The condition $p\gtrsim N^{\beta}$ is technical and serves to quantify the difficulty of distinguishing the least favorable parameters by calculating the divergences between carefully constructed distributions. Despite the extensive literature on integrative precision matrix estimation as reviewed in Section 1.1, to the best of our knowledge, Theorems 3-4 are among the first to establish the minimax theories. The lower bound techniques and the matrix constructions in the proofs of Theorems 3-4 are novel and are of independent interest to the integrative precision matrix analysis under heterogeneity.

### v2

@@ 4 Theoretical Properties
where $r\geq 1$ is a fixed constant, and the weighted $\ell_{1}$ norm $\|\cdot\|_{1,w}$ and weighted $\ell_{2}$ norm $\|\cdot\|_{2,w}$ for a vector $a\in\mathbb{R}^{M}$ are given by
$\|a\|_{1,w}:=\frac{1}{N}\sum_{m=1}^{M}n_{m}|a_{m}|\text{ and }\|a\|_{2,w}:=\sqrt{\frac{1}{N}\sum_{m=1}^{M}n_{m}a_{m}^{2}}.$
Therefore, (17)-(18) quantify the entry-wise integrative deviation between $A^{(\cdot)}$ and $B^{(\cdot)}$ under the $r$-th power of the weighted $\ell_{1}$ and $\ell_{2}$ norms, respectively. It is noteworthy that the sample size heterogeneity across the local sites is accommodated by the weights $\{n_{m}/N\}_{m\in[M]}$, and consequently there is no necessity to impose the commonly assumed balanced sample size condition to derive the convergence rates.
@@ 4.1 Rates of Convergence for HEAT
The thresholding function $T_{1,\lambda}:\mathbb{R}\rightarrow\mathbb{R}$ satisfies the properties (15), and the function $T_{2,\lambda}:\mathbb{R}^{M}\rightarrow\mathbb{R}^{M}$ satisfies the properties (16) with norm $\|\cdot\|=\|\cdot\|_{2,w}$ on $\mathbb{R}^{M}$.
@@
Assumptions 1-3 impose mild conditions on data distributions, and Assumption 4 summarizes the requirements for the thresholding functions. Specifically, Assumption 1 is a regularity condition to prevent singular covariances and avoid diverging variances. Assumption 2 excludes the heavy-tailed distributions, and Assumption 3 assumes non-degenerate variances for the influence functions $\phi_{i,j,k}^{(m)}$’s. Assumptions 2-3 can be implied by Assumption 1 if the distributions are Gaussian. The weighted norm $\|\cdot\|_{2,w}$ for $T_{2,\lambda}$ in Assumption 4 accommodates the sample size differences among the distributed sites. Section LABEL:sec:addition_choice_threshold of the supplement presents specific examples of $T_{1,\lambda}$ and $T_{2,\lambda}$ that satisfy Assumption 4, including soft, hard, SCAD and MCP thresholding functions.
@@
Under Assumptions 1-3, assume $s_{0}\log(p)=o(n)$, $n\geq Cs_{0}^{2}\log(p)$ for some constant $C>0$, and $\lambda_{j}^{(m)}\asymp\sqrt{\log(p)/n_{m}}$ uniformly for $j\in[p]$ and $m\in[M]$, and define
@@
and $\tau=\widebar{C}_{3}(n_{\max}/n)M^{5/2}\log(p)/n^{1/2}$ for some constants $\widebar{C}_{2},\widebar{C}_{3}>0$, then for any constant $\widebar{\delta}>0$, the event $\bigcap_{j,k\in[p]}\left\{\|\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}-\Lambda_{j,k}^{(\cdot)}\|_{2,w}\leq\widebar{\lambda}_{2,j,k}\right\}$ holds with probability tending to 1.
Propositions 1-2 characterize the stochastic volatility of $\widebar{\Omega}_{j,k}$ and $\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}$. As a result, $|\widebar{\Omega}_{j,k}|\leq\widebar{\lambda}_{1,j,k}$ and $\|\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}\|_{2,w}\leq\widebar{\lambda}_{2,j,k}$ respectively indicate that the underlying signals $\Gamma_{j,k}$ and $\Lambda_{j,k}^{(\cdot)}$ are weak, allowing for thresholding $\widebar{\Omega}_{j,k}$ or $\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}$ to exactly zero. The variance terms $\|v_{j,k}^{(\cdot)}\|_{1,w}$, $\|v_{j,k}^{(\cdot)}\|_{1},\|v_{j,k}^{(\cdot)}\|_{2},\|v_{j,k}^{(\cdot)}\|_{\infty}$ in $\widebar{\lambda}_{1,j,k}$ and $\widebar{\lambda}_{2,j,k}$ capture entry-dependent volatility. The remaining terms, $s_{0}M\log(p)/N$ in Proposition 1 and $s_{0}\sqrt{M}\log(p)/\sqrt{nN}$ in Proposition 2, represent higher-order errors of the local estimators (6).
The additional term $\tau/\sqrt{N}$ in Proposition 2 arises from the Chi-square approximation applied to the weighted sum-of-square quantity $\|\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}-\Lambda_{j,k}^{(\cdot)}\|_{2,w}$ as presented in Lemma LABEL:lemma:ss_chis_bound of the supplement. It is negligible if $\tau^{2}=o(M\vee\log(p))$, and in the subsequent analysis this condition is assumed for simplification and clarity. We note that this is a technical assumption that may be relaxed with further theoretical refinement; see Section LABEL:sec:more_discuss_chi of the supplement for more detailed discussions. A key advantage of our methodology is that, unlike typical concentration results for quadratic forms or polynomials, the Chi-square approximation developed here provides both explicit constant and tight convergence rate, which facilitates both practical implementation and theoretical error analysis.
In practice, as stated in Section 3, the determination of shrinkage levels depends on whether the local variance estimators $(\widehat{v}_{j,k}^{(m)})_{j,k,m}$ are transmitted from the local sites to the central site. For the case where the variance estimators are available at the central site and are used to construct the entry-dependent thresholds based on Propositions 1-2, the following Proposition 3 establishes the uniform rate of convergence for the variance estimation.
@@
Based on the results in Propositions 1-3, the shrinkage levels $\lambda_{1,j,k}$ and $\lambda_{2,j,k}$ in HEAT steps (13)-(14) can be specified either universally, applying a uniform threshold across all entries:
$\lambda_{1,j,k}=$ $\lambda_{1}\asymp\sqrt{\frac{\log(p)}{N}}+\frac{s_{0}M\log(p)}{N},$ (19)
$\lambda_{2,j,k}=$ $\lambda_{2}\asymp\sqrt{\frac{M+\log(p)}{N}}+\frac{\sqrt{M}s_{0}\log(p)}{\sqrt{nN}},$
or adaptively using entry-dependent thresholds that incorporate local variance estimates:
@@
$\lambda_{2,j,k}=$ $(1+\delta)\sqrt{\frac{\|\widehat{v}_{j,k}^{(\cdot)}\|_{1}+2\sqrt{2}\|\widehat{v}_{j,k}^{(\cdot)}\|_{2}\sqrt{\log(p)}+4\|\widehat{v}_{j,k}^{(\cdot)}\|_{\infty}\log(p)}{N}}+\widebar{C}_{2}\frac{\sqrt{M}s_{0}\log(p)}{\sqrt{nN}},$
where $\delta>0$ is a small constant. With these two specifications for the shrinkage levels, the subsequent theorems establish the rates of convergence for the HEAT estimator $\widetilde{\Omega}^{(\cdot)}$ in (12) under the loss matrices $L_{1,r}$ and $L_{2,r}$ in (17)-(18). Specifically, Theorem 1 evaluates the integrative estimation losses under the general matrix $l$-norm with $l\in[1,\infty]$, while Theorem 2 investigates the matrix Frobenius norm, which essentially vectorizes the loss matrices and measures the accuracy through vector $\ell_{2}$ norm.
@@
Under Assumptions 1-4, if $s_{0}\log(p)=o(n)$, $\log^{3}(p)=o(N)$, $n\geq Cs_{0}^{2}\log(p)$ for some constant $C>0$, $\tau^{2}=o(M\vee\log(p))$ and $\lambda_{j}^{(m)}\asymp\sqrt{\log(p)/n_{m}}$ uniformly for $j\in[p]$ and $m\in[M]$, and additionally assume $\log^{2}(p\vee n_{\max})(s_{0}\vee\log(p))=o(n)$ for the shrinkage levels in (20), then for any $r\geq 1$ and $l\in[1,\infty]$, we have
@@
The convergence rates in Theorem 1 consist of two components: $s_{1}(\varphi_{1})^{r}$ and $s_{2}(\varphi_{2})^{r}$, which respectively reflect the estimation errors under the homogeneity and the heterogeneity. The same decomposition and interpretation also apply to Theorem 2. Note that both $\varphi_{1}$ and $\varphi_{2}$ consist of two terms. The terms $\sqrt{\frac{\log(p)}{N}}$ and $\sqrt{\frac{M+\log(p)}{N}}$ are the leading statistical errors for integration, while $\frac{s_{0}M\log(p)}{N}$ and $\frac{\sqrt{M}s_{0}\log(p)}{\sqrt{nN}}$ are the higher-order errors induced by individual estimation, representing the additional costs of distributed data aggregation. Consequently, the higher-order errors will be dominated by the leading error terms, i.e.,
$\varphi_{1}\asymp\psi_{1}:=\sqrt{\frac{\log(p)}{N}}\text{ and }\varphi_{2}\asymp\psi_{2}:=\sqrt{\frac{M+\log(p)}{N}},$ (21)
if $\frac{M^{2}}{N}\lesssim{(s_{0}^{2}\log(p))}^{-1}$ and $s_{0}^{2}\log^{2}(p)\lesssim(1\vee\frac{\log(p)}{M})n$ hold, which are further guaranteed by $\frac{M}{n}\lesssim(s_{0}^{2}\log(p))^{-1}$. These restrictions imply that the number of distributed sites can not be too large, and similar requirements also appear in the distributed algorithms under homogeneity (e.g., Zhang et al.,, 2015; Lee et al.,, 2017; Battey et al.,, 2018; Jordan et al.,, 2019) and heterogeneity (e.g., Maity et al.,, 2022; Cai et al.,, 2022). In the following section, it will be shown that $\psi_{1}$ and $\psi_{2}$ in (21) achieve the minimax optimal rates.
@@ 4.2 Minimax Lower Bounds for Integrative Estimation
Suppose that $X_{i}^{(m)}\sim N(\mu^{(m)},\Sigma^{(m)})$ for $i\in[n_{m}],m\in[M]$ and Assumption 1 holds. If $3\leq s_{1}<p/2$, $3\leq s_{2}<p/2$, $s_{0}^{2}\log^{3}(p)=o(N)$, $s_{2}^{2}M=o(N)$, $N\lesssim nM$ and $p\gtrsim N^{\beta}$ for some $\beta>1$, then for any $r\geq 1$ and $l\in[1,\infty]$,
$\inf_{\widetilde{\Omega}^{(\cdot)}}\sup_{\Omega^{(\cdot)}\in\Theta}\mathbb{E}\|L_{1,r}(\widetilde{\Omega}^{(\cdot)},\Omega^{(\cdot)})\|_{l}\gtrsim s_{1}(\psi_{1})^{r}+s_{2}(\psi_{2})^{r},$
$\inf_{\widetilde{\Omega}^{(\cdot)}}\sup_{\Omega^{(\cdot)}\in\Theta}\mathbb{E}\|L_{2,r}(\widetilde{\Omega}^{(\cdot)},\Omega^{(\cdot)})\|_{l}\gtrsim s_{1}(\psi_{1})^{r}+s_{2}(\psi_{2})^{r}.$
@@
$\inf_{\widetilde{\Omega}^{(\cdot)}}\sup_{\Omega^{(\cdot)}\in\Theta}\frac{1}{p}\mathbb{E}\|L_{1,r}(\widetilde{\Omega}^{(\cdot)},\Omega^{(\cdot)})\|_{\text{F}}^{2}\gtrsim s_{1}(\psi_{1})^{2r}+s_{2}(\psi_{2})^{2r},$
$\inf_{\widetilde{\Omega}^{(\cdot)}}\sup_{\Omega^{(\cdot)}\in\Theta}\frac{1}{p}\mathbb{E}\|L_{2,r}(\widetilde{\Omega}^{(\cdot)},\Omega^{(\cdot)})\|_{\text{F}}^{2}\gtrsim s_{1}(\psi_{1})^{2r}+s_{2}(\psi_{2})^{2r}.$
Theorems 3-4 establish the minimax lower bounds under the matrix $l$-norm and Frobenius norm, respectively. The condition $p\gtrsim N^{\beta}$ is technical and serves to quantify the difficulty of distinguishing the least favorable parameters by calculating the divergences between carefully constructed distributions. Conditions $s_{1}\geq 3$ and $s_{2}\geq 3$ are also technical in constructing the least favorable distributions, and the results in Theorems 3-4 remain valid when either $s_{1}$ or $s_{2}$ equals 0, as detailed in the proofs. Despite the extensive literature on integrative precision matrix estimation as reviewed in Section 1.1, to the best of our knowledge, Theorems 3-4 are among the first to establish the minimax theories. The lower bound techniques and the matrix constructions in the proofs of Theorems 3-4 are novel and are of independent interest to the integrative precision matrix analysis under heterogeneity.